Demand for C&A Resort has a seasonal pattern. Total demand next year is expected to be 1,525. Given the historical data of last two years, what is the seasonal index for summer? Season Year 1 Year 2 Fall Winter156 212 Spring 489 518 Summer314 352 352 391 O 1.449 1.068 365 0.957 0.526
Answer #1

Answer: Option D:0.957

Average for demand for year 2 =(391+212+518+352)/4=368.25

Seasonal index for summer =352/368.35=0.956